if i have a acid solution at 5 % concentration and another one at 10% concentration how much of each solution do I need to obtain 50ml at 9% concentration
Answer by Alan3354(69443)

if i have a acid solution at 5 % concentration and another one at 10% concentration how much of each solution do I need to obtain 50ml at 9% concentration
-----------------------
f = amount of 5%
t = amount of 10%
-----
f + t = 50 (total solution)
5f + 10t = 50*9 (% acid)